# Small Batch Size Training for Language Models: When Vanilla SGD Works, and Why Gradient Accumulation Is Wasteful

### v1

@@ A.1 Model architecture
Across all of our pretraining experiments, we used four decoder-only transformer models [77] that share the same architecture and only different in layer dimensions and tokenizer. We specify the exact layer dimensions and tokenizers used in Table 1.
We used a model architecture based on GPT-2 [59] with several small changes based on modern practices to improve model FLOPs utilization (MFU) [76], simplify implementation, and marginally improve model performance:
@@
    • We increased the head dimension on the smaller models from 64 to 128 to achieve a higher MFU. We trained each model on TPU v4, which has 128 x 128 multiply-accumulators. Using a head dimension any lower than 128 would prevent us from filling the multiply-accumulators, thereby lowering MFU. This decision follows the Gemma family of models optimized for TPU training [69].
  • RMSNorm [63]
@@
    • We do not tie the weights of the first and last layer. This increases number of trainable parameters but does not increase the number of active parameters or the cost of training.
  • Rotary positional embeddings (RoPE) [61]
  • QK-norm [62]
  • GELU [64]
We trained most models to follow Chinchilla scaling laws, i.e. using 20 tokens per active parameter. The only exception to this rule is the GPT-3 (1.3B) model, which we could only afford to train for 10B tokens, similar to the original GPT-2 (1.5B) model.
Model 30M 19M[15] GPT-2 (124M) GPT-3 (1.3B)
@@
A.2 Computational cost
We used a pod of 32 TPU v4 chips to train all models [78]. Table 2 shows the time for a single training run of each model as well as the total computational cost for each model. Using batch sizes 1 and 2 resulted in roughly 30-70% drop in MFU compared to using a larger batch size, depending on the model size and sequence length. In practice we do not recommend using a batch size so small that it severely degrades MFU. We only used the smallest batch sizes to scientifically study the effect of batch size.
Model 30M 19M[15] GPT-2 (124M) GPT-3 (1.3B) Gemma3 (4B)
@@ A.2 Computational cost
Num. runs: Footnote 2 3110 / 312 11 /
Num. runs: Figure 3 1911 / / / /
Num. runs: Figure 4 1820 / / / /
Num. runs: Figure 5 / 1260 / / /
Num. runs: Figure 6 / / 40 / /
Num. runs: Figure 7 / / / / 256
@@ Appendix B Additional Results
In Figure 9 we illustrate on a toy problem why momentum is required when the batch size is large but not required when the batch size is small. We run SGD on two variables $(x,y)$ to minimize the value of a loss function defined as $x+10y^{2}$. Notice that this loss function is much steeper in the vertical direction compared to the horizontal direction.
We simulate the use of a small and a large batch size by adding noise to the gradient estimate at every step. In particular, we define the minibatch gradient estimate at each step to be the the true gradient multiplied by a random variable sampled from a normal distribution $\mathcal{N}(1,\sigma^{2})$, where the noise scale $\sigma$ controls the signal-to-noise ratio of the minibatch gradient estimate. We simulate a large batch size by using a high signal-to-ratio of 5, and a small batch size by using a small signal-to-noise ratio of only 0.3.
In our experiments training language models, we fixed the computational budget across all batch sizes by only training for a single epoch. Analogously in this toy example, we only run 10 optimization steps using the simulated large batch size, but we run 100 optimization steps using the small batch size. The idea is that if we use a small batch size, the minibatch gradient estimates at each step become more noisy but in turn we get to take more optimizer steps.
@@
In the top row of Figure 9, we see that when a large batch size is used, we only get to take a small number of steps, and so in turn every step has to be large. Because the loss function is much steeper in the vertical direction compared to the horizontal direction, the large step size results in oscillations along the vertical direction. In this case, using momentum helps damped the oscillations along the vertical direction and speeds up convergence along the horizontal direction. In fact, in the extreme case of full-batch gradient descent (which has no gradient noise), using momentum probably improves the convergence rate of gradient descent for this toy problem [65].
We show the small batch size setting in the bottom row of Figure 9. Notice that since the batch size is small, every step becomes more noisy, but in turn we can take more steps and reduce the size of each step. As a result of decreasing the step size, the optimizer no longer overshoots along the vertical direction, obviating the need for momentum. There are no longer any oscillations to dampen, so using SGD with and without momentum results in similar performance.

### v2

@@ A.1 Model architecture
Across all of our pretraining experiments, we used four decoder-only transformer models [Vaswani et al., 2017] that share the same architecture and only differ in layer dimensions and tokenizer. We specify the exact layer dimensions and tokenizers used in Table˜1.
We used a model architecture based on GPT-2 [Radford et al., 2019] with several small changes based on modern practices to improve model FLOPs utilization (MFU) Chowdhery et al. [2023], simplify implementation, and marginally improve model performance:
@@
    • We increased the head dimension on the smaller models from 64 to 128 to achieve a higher MFU. We trained each model on TPU v4, which has 128 x 128 multiply-accumulators. Using a head dimension any lower than 128 would prevent us from filling the multiply-accumulators, thereby lowering MFU. This decision follows the Gemma family of models optimized for TPU training [Team et al., 2025].
  • RMSNorm [Zhang and Sennrich, 2019]
@@
    • We do not tie the weights of the first and last layer. This increases the number of trainable parameters but does not increase the number of active parameters or the cost of training.
  • Rotary positional embeddings (RoPE) [Su et al., 2024]
  • QK-norm Dehghani et al. [2023]
  • GELU [Hendrycks and Gimpel, 2016]
We pretrained most models to follow Chinchilla scaling laws, i.e. using 20 tokens per active parameter. The only exception to this rule is the GPT-3 (1.3B) model, which we only trained for 10B tokens, similar to the original GPT-2 (1.5B) model.
Model 30M 19M[Xiao, 2024] GPT-2 (124M) GPT-3 (1.3B)
@@
A.2 Learning rate schedule
For pretraining, we used a linear warmup for the first 5% of the training steps (from zero to peak learning rate), followed by cosine decay (from peak learning rate to zero). We use the same schedule for all batch sizes. For fine-tuning, we used a constant learning rate.
Figure 9: Stochastic rounding enables bfloat16 training. We repeat the training runs from Figure 1a using the optimum hyperparameters found for Adafactor. We ablate the precision of the model weights, using: (1) float32 – same as in Figure 1a, (2) bfloat16 with closest rounding, (3) bfloat16 with stochastic rounding. The optimizer state is always stored in float32 precision since it has a negligible size.
A.3 Stochastic rounding
Across all experiments, we compute activations and store the optimizer state in float32 precision. We also store all trainable model parameters in float32 precision except for Adafactor in Figure˜8, which uses bfloat16 weights in order to match the memory footprint of LoRA. While it is a common practice to use bfloat16 weights for computation, the master copy of model weights is typically still stored in float32 to enable accumulating small updates to the weights. When the master weights are stored in bfloat16 (which only has $\sim$2.4 decimal points of precision), standard deterministic rounding can bias the weight updates, meaning that accumulating gradients across many small batches no longer approximates computing gradients from a single larger batch. Because of this bias, reducing the batch size might actually be detrimental with standard deterministic rounding. One approach to remove this bias is to compute updated weights in float32, then stochastically round them to bfloat16 for storage. In a pretraining experiment in Figure˜9, we show that stochastic rounding can bring the performance of bfloat16 weights close to the performance of float32 weights, even for the smallest batch sizes. In the fine-tuning experiment in Figure˜8, we did not observe a statistically significant difference between float32 weights and bfloat16 weights with stochastic rounding for Adafactor.
A.4 Computational cost
We used a pod of 32 TPU v4 chips to train all models [Jouppi et al., 2023]. Table˜2 shows the time for a single training run of each model as well as the total computational cost for each model. Using batch sizes 1 and 2 resulted in roughly a 30-70% drop in MFU compared to using a larger batch size, depending on the model size and sequence length. In practice, we do not recommend using a batch size so small that it severely degrades MFU. We only used the smallest batch sizes to scientifically study the effect of batch size.
Model 30M 19M[Xiao, 2024] GPT-2 (124M) GPT-3 (1.3B) Gemma3 (4B)
@@ A.4 Computational cost
Num. runs: Figure˜1 3110 / 312 11 /
Num. runs: Figure˜3 1911 / / / /
Num. runs: Figure˜4 1820 / / / /
Num. runs: Figure˜5 / 1260 / / /
Num. runs: Figure˜6 / / 40 / /
Num. runs: Figure˜8 / / / / 256
@@ Appendix B Additional Results
In Figure˜10 we illustrate on a toy problem why momentum is required when the batch size is large but not required when the batch size is small. We run SGD on two variables $(x,y)$ to minimize the value of a loss function defined as $x+10y^{2}$. Notice that this loss function is much steeper in the vertical direction compared to the horizontal direction.
We simulate the use of a small and a large batch size by adding noise to the gradient estimate at every step. In particular, we define the minibatch gradient estimate at each step to be the true gradient multiplied by a random variable sampled from a normal distribution $\mathcal{N}(1,\sigma^{2})$, where the noise scale $\sigma$ controls the signal-to-noise ratio of the minibatch gradient estimate. We simulate a large batch size by using a high signal-to-ratio of 5, and a small batch size by using a small signal-to-noise ratio of only 0.3.
In our experiments training language models, we fixed the computational budget across all batch sizes by only training for a single epoch. Analogously in this toy example, we only run 10 optimization steps using the simulated large batch size, but we run 100 optimization steps using the small batch size. The idea is that if we use a small batch size, the minibatch gradient estimates at each step become more noisy, but in turn we get to take more optimizer steps.
@@
In the top row of Figure˜10, we see that when a large batch size is used, we only get to take a small number of steps, and so in turn every step has to be large. Because the loss function is much steeper in the vertical direction compared to the horizontal direction, the large step size results in oscillations along the vertical direction. In this case, using momentum helps dampen the oscillations along the vertical direction and speeds up convergence along the horizontal direction. In fact, in the extreme case of full-batch gradient descent (which has no gradient noise), using momentum probably improves the convergence rate of gradient descent for this toy problem [Goh, 2017].
We show the small batch size setting in the bottom row of Figure˜10. Notice that since the batch size is small, every step becomes more noisy, but in turn we can take more steps and reduce the size of each step. As a result of decreasing the step size, the optimizer no longer overshoots along the vertical direction, obviating the need for momentum. There are no longer any oscillations to dampen, so using SGD with and without momentum results in similar performance.